OpenStudy (anonymous):

Snookers lumber can convert logs into either lumber or plywood. In a given day, the mill turns out three times as many units of plywood as lumber. It makes a profit of 30 on a unit of lumber and 45 on a unit of plywood. How many of each unit must be produced and sold to make a profit of $16665?

OpenStudy (saifoo.khan):

Snookers Lumber can convert logs into either lumber or plywood. In a given day,the mill turns out three times as many units of plywood as lumber. It makes a profit of $30 on a Unit of lumber and $45 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of 16665.?: Let L = no. of units of lumber Let P = no. of units of plywood : "In a given day,the mill turns out three times as many units of plywood as lumber." P = 3L : "It makes a profit of $30 on a Unit of lumber and $45 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of 16665.?" 30L + 45P = 16665 Replace P with 3L 30L + 45(3L) = 16665 30L + 135L = 16665 165L = 16665 L = L = 101 units of lumber then 3(101) = 303 units of plywood